### v2

@@ B.1.2 Task Significance
Legal case summarization is a foundational task for improving access to justice, supporting legal research, and facilitating informed decision-making across the legal system. Many real-world cases involve large volumes of interrelated documents. These include complaints, motions, rulings, and settlement agreements, which must be synthesized into concise, coherent summaries. Automatic summarization addresses this issue by condensing lengthy texts into concise summaries, enabling users to quickly grasp essential details without reading the entire document [5]. This automation enhances accessibility by providing clear summaries for lawyers, scholars, and the general public, improves efficiency by assisting legal professionals in quickly understanding case details, and supports research by highlighting crucial aspects of cases. By reducing the time and effort required to comprehend extensive legal materials from legal experts, automatically generated summaries streamlines legal workflows and aids in informed decision-making. This challenge becomes even more pronounced in domains where litigation is prolonged, procedurally complex, and socially impactful.
@@
One such domain is civil rights litigation, which has played a pivotal role in reforming public institutions and shaping policy in areas such as education, policing, incarceration, and disability rights. These cases often span years and generate hundreds of filings, yet most remain inaccessible to the broader public, scholars, and even practitioners. The Civil Rights Litigation Clearinghouse [^4] https://clearinghouse.net/, a leading repository in this space, curates extensive documentation from landmark and ongoing civil rights cases—many of which have never resulted in published judicial opinions. As a result, key legal developments are often buried in unstructured documents that are difficult to access and interpret without expert knowledge. Multi-document case summarization helps bridge the accessibility gap. It enables non-experts to understand legal developments, allows legal professionals to quickly grasp the structure and outcome of complex cases, supports advocacy and policymaking by surfacing systemic patterns, and reduces the burden of reviewing lengthy legal documents, giving researchers easier access to representative case content. We collect data from Clearinghouse, with further details provided in Section B.1.3.
@@
A related work, Multi-LexSum [5], also investigates the task of multi-document legal case summarization in civil rights litigation domain using data from Clearinghouse. However, their study relies on automatic evaluation metrics such as ROUGE-{1,2,L} [16] and BERTScore [18]. In contrast, our work constructs a new dataset by collecting cases data from the Clearinghouse and introducing clear selection criteria with legal researchers grounded in both diversity and difficulty, where diversity refers to the range of case topics and difficulty is based on structural factors such as such as the number of documents, length, and complexity of legal proceedings. Furthermore, we develop a fine-grained, checklist-based evaluation rubric in collaboration with a legal researcher to ensure comprehensive coverage of key case elements. To assess model performance more reliably, we design a structured evaluation pipeline leveraging LLMs as evaluators, enabling nuanced analysis beyond surface-level lexical similarity.
@@ B.1.3 Data Acquisition and Preprocessing
To further ensure the quality of the gold-standard case summary, we discuss with legal researchers and identify the required checklist items that must be included in the summary. We then use GPT-4o as a judge to assess whether the human-written reference covers the information and select the cases with highest-quality human summaries, determined by the highest coverage of checklist items. The specific prompt used for accessing the reliability of human-written summary is shown in Tab. 3. Based on the judgment results, we finally select 100 samples with high-quality and long human reference summaries.
@@ B.1.4 Illustrative Example
An illustrative example is shown in Tab. 4. All samples for T1 are publicly available [^7] https://huggingface.co/datasets/launch/ExpertLongBench.
@@
Additionally, Tab. 5 presents the model prompt. The prompt was carefully constructed to be high-level and general instead of providing the exact checklist items. The inputs for this prompt are the documents shown in Tab. 4.
@@ B.1.6 Checklist-mapped Reference
For the human reference, we used GPT-4o to extract answers to each checklist item individually. Each prompt targeted a single item and followed a structured format. To improve contextual grounding, we used role-playing to frame the model as an assistant to a lawyer. If no related information was mentioned in the legal case summary, the model should return “N/A” to reduce the possibility of hallucination. An example is shown in Tab. 6.
For the model output, we use Qwen2.5-72B to create the checklist-mapped reference. Our decision to use a different model is supported in §4.1. To reduce cost while maintaining quality, we grouped the checklist items rather than prompting for them individually. We divide the 26 items into five groups, each with four to six checklist items, based on the average expected length of the model’s response. Each model output prompt used the same template, role-playing format as the human reference prompt. Due to space constraints, we release the full prompt set in our public GitHub repository.
The checklist-mapped reference of the example in Tab. 4 is presented in Tab. 7. To assess the quality of the checklist-mapped reference, we selected an additional 30 difficult and diverse samples, different from the main dataset, and conducted human verification to examine the faithfulness of the model response. Details of this experiment can be found in LABEL:ap:validation-of-checklist.
@@ B.2.1 Task Definition
A Statement of Fact (SOF) is a document used in legal cases to persuade the judge of a certain viewpoint. It outlines the events and circumstances leading up to a legal dispute in an objective manner, while also incorporating persuasive language when appropriate [39]. The objective is to generate a comprehensive account of the case, based on courtroom transcripts, that accurately reflects the its development. The input includes transcripts from multiple proceedings, such as preliminary trials, adjudication hearings, and termination trials. The output should present all relevant facts, direct quotations, and procedural details with proper attribution and clear alignment to legal standards. It must stand alone as a complete summary, without requiring reference to the original transcripts.
@@ B.2.2 Task Significance
Traditionally, drafting such briefs is a time-intensive and expert-level task. Attorneys must review hundreds of pages of transcripts to extract and organize key information into create an coherent narrative. This process often takes 20 to 40 hours or more, even for seasoned professionals [40]. Manual note-taking adds further difficulty - often leading to inconsistent phrasing, missed information or subjective interpretation. It also requires frequent references to transcripts, which not only slows progress, but increases the risk of delays and oversights in case processing.
@@ B.2.3 Data Acquisition and Preprocessing
For .pdf files, we first attempted direct text extraction using PyPDF2888https://pypi.org/project/PyPDF2/. If the PDF lacked embedded text (e.g., scanned images), we applied OCR using OCRmyPDF [^9] https://github.com/ocrmypdf/OCRmyPDF with deskewing, noise cleaning, and compression optimizations. This was the library recommended to us by our collaborators.
@@
To filter 113 transcript samples, we selected the 100 with the longest human-written reference SOFs, consistent with the reasoning in §3.2. On average, the 100 samples included 14 transcript documents, each 5K+ tokens in length, and paired with a human reference SOF also exceeding 5K tokens. This filtering ensured that our dataset contained the most information-dense and well-documented cases.
@@ B.2.4 Illustrative Example
Due to the proprietary nature of the data, we cannot provide a specific example. [^14] For review purpose, data can be previewed via the dataset url in the submission form. However, in Tab. 8, we provide the model prompt for obtaining model outputs.
@@ B.2.6 Checklist-mapped Reference
The creation of the checklist-mapped reference is for T2 is similar to T1’s process in described in Section B.1.6.
However, unlike in T1, the checklist items used to construct the model output checklist-mapped reference are grouped differently. We divide the 41 items into nine groups based on: (1) the average expected length of the model’s response, and (2) the topical category of the item (e.g., Details of service plan, Impact of each court event, etc.). For instance, we grouped checklist items 1–6 together, as they all fall under the categories Initial incident and Background situation. This grouping strategy helps the model process large inputs more effectively by allowing it to focus on fewer sections at a time. The prompt structure for extracting an individual checklist item is illustrated in Tab. 9.
@@ B.3.2 Task Significance
A material can often be synthesized through multiple methods, and the choice of precursors and synthesis conditions can significantly affect the properties and yields of the resulting material [41]. Understanding the reasoning behind these choices is essential for materials scientists aiming to optimize synthesis protocols and develop new materials with desirable properties through more efficient processes. Additionally, materials discovery efforts, especially those aided by high-throughput computations [42, 43], can be accelerated by prioritizing synthesis recipes that are grounded in the most promising mechanistic insights. Per our discussion with materials science researchers, writing the explanations for a synthesis recipe from scratch would take a well-trained materials science PhD student 1-2 hours.
@@ B.3.3 Data Acquisition and Preprocessing
Despite the availability of large-scale synthesis recipe datasets [43, 44], these curated recipes are not accompanied by explanations. Even for well-trained materials science PhDs, writing explanations for synthesis recipes from scratch is challenging and time-consuming. In contrast, given statements about the synthesis recipe, undergraduate-level chemistry students are able to distinguish between explanations and non-explanations. Therefore, we adopt a semi-automated approach to collect synthesis recipes along with corresponding explanations from existing literature.
@@
We retrieve papers containing the keyword “solid-state synthesis” from Science [^17] https://www.science.org and Wiley [^18] https://onlinelibrary.wiley.com, as the aspects of interest are more relevant to solid-state synthesis. The text of the paper is extracted directly from the HTML version of the paper. During the automated extraction step, we generate five different outputs using GPT-4o for each paper to increase the recall of possible explanations. To improve the faithfulness of the extraction, we also prompt the model to provide the source sentences corresponding to each explanation. The prompt is shown in Tab. 10, which is based on the rubric designed in Section B.3.5. Feedback from the annotators indicates that some explanations extracted by the LLMs describe properties of the synthesized materials (e.g., crystal structure, conductivity) rather than justifying the synthesis steps themselves.
@@ B.3.4 Illustrative Example
An illustrative example is shown in Tab. 11. All samples for T3 are publicly available.
@@
Additionally, Tab. 12 presents the model prompt.
@@
You are a materials science researcher. Given a synthesis recipe that includes the target material, selected precursors, and synthesis steps, your task is to justify the key decisions made in the recipe. This includes explaining the rationale behind the choice of precursors, reaction conditions, and processing steps, using relevant principles such as structural compatibility, chemical reactivity, and desired phase formation. Output the explanation rationales as a list of bullet points, where each bullet point contains complete sentences.
@@ B.3.6 Checklist-mapped Reference
Synthesis explanation generation does not involve the checklist mapping process, as the annotation process directly produces the reference checklist. An example checklist-mapped reference is shown in the illustrative example in Tab. 11.
@@ B.4.1 Task Definition
Assessing pedagogical alignment involves evaluating how well a large language model (LLM) can replicate effective teaching strategies, such as offering scaffolded guidance without directly revealing the answer [45, 46, 47, 1]. One possible approach is to evaluate LLMs through real-world interactions with students. However, this poses challenges, as student responses can vary unpredictably across different models [48]. Additionally, using the same group of students to test multiple LLMs introduces complications: once students are exposed to a topic in one session, their familiarity with the material can influence their performance in subsequent sessions, making it difficult to isolate the effectiveness of each system. Conversely, using different students for the same set of topics may lead to non-standardized assessments due to variations in their prior knowledge and skill levels. The objective is to assess whether LLMs can provide pedagogically sound feedback that supports student learning without directly giving away answers.
To provide a repeatable and standardized framework for evaluating the pedagogical abilities of different LLMs, we propose a task focused on assessing the effectiveness and pedagogical alignment of feedback within a tutor-student dialogue. Each task instance presents a multi-turn conversation in which a student attempts to solve a biology problem with the guidance of a tutor. The tutor employs a step-by-step problem-solving strategy, breaking the larger problem into smaller sub-problems and promoting active learning by guiding the student through each segment [46]. This process is sequential: the student addresses one sub-problem at a time before moving on to the next. The conversational context includes the full dialogue history up to the current sub-problem, concluding with the student’s latest response.
@@ B.4.2 Task Significance
First, LLMs may produce inaccurate reasoning [49] or rely on incorrect information [50] when generating feedback, which can hinder rather than support student learning. Second, while LLMs are generally optimized to be helpful and harmless, the operational definition of “helpfulness” may not align with pedagogical goals. Specifically, a fundamental strategy in conceptual learning is to guide students through problem-solving processes without directly revealing the answer. In contrast, LLMs are often trained to prioritize immediate assistance, which may conflict with the educational objective of fostering deep understanding through indirect guidance.
@@ B.4.3 Data Acquisition and Preprocessing
We first provide a background on the source of the data that will be used for building this task. We used the multi-turn student-tutor conversations simulated using GPT-4o by [51]. These dialogues are grounded in the socio-constructivist model of learning [52], where the tutor adopts a supportive and encouraging tone while offering feedback through Socratic questioning and indirect hints, rather than direct instruction. An illustrative example of such a conversation is provided in Tab. 15. Each simulated interaction features a student working through a complex biology problem that requires multi-step reasoning and problem decomposition, with the tutor guiding the student through the process, making the dataset well-suited for our evaluation. Human annotators validated each simulated example to ensure factual accuracy and alignment with the socio-constructivist learning model.
@@
In contrast, we concentrate solely on the non-affective components of the feedback, as prior studies have shown minimal influence of positive language and praise on student outcomes [53, 54]. Our primary objective is to develop evaluation criteria that assess the effectiveness of feedback in promoting conceptual understanding. Studies have consistently found that feedback elements which identify problems and errors and propose solutions are the most beneficial for student learning [55, 56, 57, 58]. Accordingly, our task is designed to evaluate an LLM’s ability to generate feedback that provides accurate problem diagnosis and rectification. While indirect feedback—such as hints or questions that avoid explicitly identifying the mistake or providing the answer—can foster meta-cognitive and conceptual learning, its effectiveness may diminish when students repeatedly struggle with the same sub-problem [59]. Accordingly, our evaluation also emphasizes the LLM’s ability to generate contextually appropriate feedback based on the nature of the student’s mistakes for a given sub-problem. In cases where the student makes repeated errors on the same sub-problem, the LLM should accurately recognize this pattern and respond with explicit error identification and corrective guidance. In this section, we explain how to get fine-grained information from the seed data collected by [51].
@@
Consider a single response of the tutor from the example shown in Tab. 15. We focus on the second response from the tutor and provide fine-grained information associated with it in Tab. 13.
@@
To assess effectiveness in error identification and rectification, we extract these components from the feedback generated by the simulated tutor. This extraction is carried out using GPT-4o for automated annotation, with subsequent human verification to ensure the accuracy of the identified issues and corresponding corrective suggestions. Since the task does not require deep domain-specific knowledge, the manual verification was performed by one of the authors with strong proficiency in English. An illustrative example of the extracted elements from a single data point is provided in Tab. 14.

### v3

@@ B.1.2 Task Significance
Legal case summarization is a foundational task for improving access to justice, supporting legal research, and facilitating informed decision-making across the legal system. Many real-world cases involve large volumes of interrelated documents. These include complaints, motions, rulings, and settlement agreements, which must be synthesized into concise, coherent summaries. Automatic summarization addresses this issue by condensing lengthy texts into concise summaries, enabling users to quickly grasp essential details without reading the entire document [Shen et al., 2022]. This automation enhances accessibility by providing clear summaries for lawyers, scholars, and the general public, improves efficiency by assisting legal professionals in quickly understanding case details, and supports research by highlighting crucial aspects of cases. By reducing the time and effort required to comprehend extensive legal materials from legal experts, automatically generated summaries streamlines legal workflows and aids in informed decision-making. This challenge becomes even more pronounced in domains where litigation is prolonged, procedurally complex, and socially impactful.
@@
One such domain is civil rights litigation, which has played a pivotal role in reforming public institutions and shaping policy in areas such as education, policing, incarceration, and disability rights. These cases often span years and generate hundreds of filings, yet most remain inaccessible to the broader public, scholars, and even practitioners. The Civil Rights Litigation Clearinghouse [^6] https://clearinghouse.net/, a leading repository in this space, curates extensive documentation from landmark and ongoing civil rights cases—many of which have never resulted in published judicial opinions. As a result, key legal developments are often buried in unstructured documents that are difficult to access and interpret without expert knowledge. Multi-document case summarization helps bridge the accessibility gap. It enables non-experts to understand legal developments, allows legal professionals to quickly grasp the structure and outcome of complex cases, supports advocacy and policymaking by surfacing systemic patterns, and reduces the burden of reviewing lengthy legal documents, giving researchers easier access to representative case content. We collect data from Clearinghouse, with further details provided in Section˜B.1.3.
@@
A related work, Multi-LexSum [Shen et al., 2022], also investigates the task of multi-document legal case summarization in civil rights litigation domain using data from Clearinghouse. However, their study relies on automatic evaluation metrics such as ROUGE-{1,2,L} [Lin, 2004] and BERTScore [Zhang et al., 2019]. In contrast, our work constructs a new dataset by collecting cases data from the Clearinghouse and introducing clear selection criteria with legal researchers grounded in both diversity and difficulty, where diversity refers to the range of case topics and difficulty is based on structural factors such as such as the number of documents, length, and complexity of legal proceedings. Furthermore, we develop a fine-grained, checklist-based evaluation rubric in collaboration with a legal researcher to ensure comprehensive coverage of key case elements. To assess model performance more reliably, we design a structured evaluation pipeline leveraging LLMs as evaluators, enabling nuanced analysis beyond surface-level lexical similarity.
@@ B.1.3 Data Acquisition and Preprocessing
To further ensure the quality of the gold-standard case summary, we discuss with legal researchers and identify the required checklist items that must be included in the summary. We then use GPT-4o as a judge to assess whether the human-written reference covers the information and select the cases with highest-quality human summaries, determined by the highest coverage of checklist items. The specific prompt used for accessing the reliability of human-written summary is shown in Tab.˜3. Based on the judgment results, we finally select 100 samples with high-quality and long human reference summaries.
@@ B.1.4 Illustrative Example
An illustrative example is shown in Tab.˜4.
@@
Additionally, Tab.˜5 presents the model prompt. The prompt was carefully constructed to be high-level and general instead of providing the exact checklist items. The inputs for this prompt are the documents shown in Tab.˜4.
@@ B.1.6 Checklist-mapped Reference
For the human reference, we used GPT-4o to extract answers to each checklist item individually. Each prompt targeted a single item and followed a structured format. To improve contextual grounding, we used role-playing to frame the model as an assistant to a lawyer. If no related information was mentioned in the legal case summary, the model should return “N/A” to reduce the possibility of hallucination. An example is shown in Tab.˜6.
For the model output, we use Qwen2.5-72B to create the checklist-mapped reference. Our decision to use a different model is supported in §4.1. To reduce cost while maintaining quality, we grouped the checklist items rather than prompting for them individually. We divide the 26 items into five groups, each with four to six checklist items, based on the average expected length of the model’s response. Each model output prompt used the same template, role-playing format as the human reference prompt. Due to space constraints, we will release the full prompt set in our public GitHub repository.
The checklist-mapped reference of the example in Tab.˜4 is presented in Tab.˜7. To assess the quality of the checklist-mapped reference, we selected an additional 30 difficult and diverse samples, different from the main dataset, and conducted human verification to examine the faithfulness of the model response. Details of this experiment can be found in Section˜E.1.
@@ B.2.1 Task Definition
A Statement of Fact (SOF) is a document used in legal cases to persuade the judge of a certain viewpoint. It outlines the events and circumstances leading up to a legal dispute in an objective manner, while also incorporating persuasive language when appropriate [LSD Law, n.d.]. The objective is to generate a comprehensive account of the case, based on courtroom transcripts, that accurately reflects the its development. The input includes transcripts from multiple proceedings, such as preliminary trials, adjudication hearings, and termination trials. The output should present all relevant facts, direct quotations, and procedural details with proper attribution and clear alignment to legal standards. It must stand alone as a complete summary, without requiring reference to the original transcripts.
@@ B.2.2 Task Significance
Traditionally, drafting such briefs is a time-intensive and expert-level task. Attorneys must review hundreds of pages of transcripts to extract and organize key information into create an coherent narrative. This process often takes 20 to 40 hours or more, even for seasoned professionals [TypeLaw, 2021]. Manual note-taking adds further difficulty - often leading to inconsistent phrasing, missed information or subjective interpretation. It also requires frequent references to transcripts, which not only slows progress, but increases the risk of delays and oversights in case processing.
@@ B.2.3 Data Acquisition and Preprocessing
For .pdf files, we first attempted direct text extraction using PyPDF2999https://pypi.org/project/PyPDF2/. If the PDF lacked embedded text (e.g., scanned images), we applied OCR using OCRmyPDF [^10] https://github.com/ocrmypdf/OCRmyPDF with deskewing, noise cleaning, and compression optimizations. This was the library recommended to us by our collaborators.
@@
To filter 113 transcript samples, we selected the 100 with the longest human-written reference SOFs, consistent with the reasoning in §3. On average, the 100 samples included 14 transcript documents, each 5K+ tokens in length, and paired with a human reference SOF also exceeding 5K tokens. This filtering ensured that our dataset contained the most information-dense and well-documented cases.
@@ B.2.4 Illustrative Example
Due to the proprietary nature of the data, we cannot provide a specific example. However, in Tab.˜8, we provide the model prompt for obtaining model outputs.
@@ B.2.6 Checklist-mapped Reference
The creation of the checklist-mapped reference is for T2 is similar to T1’s process in described in Section˜B.1.6.
However, unlike in T1, the checklist items used to construct the model output checklist-mapped reference are grouped differently. We divide the 41 items into nine groups based on: (1) the average expected length of the model’s response, and (2) the topical category of the item (e.g., Details of service plan, Impact of each court event, etc.). For instance, we grouped checklist items 1–6 together, as they all fall under the categories Initial incident and Background situation. This grouping strategy helps the model process large inputs more effectively by allowing it to focus on fewer sections at a time. The prompt structure for extracting an individual checklist item is illustrated in Tab.˜9.
@@ B.3.2 Task Significance
A material can often be synthesized through multiple methods, and the choice of precursors and synthesis conditions can significantly affect the properties and yields of the resulting material [Sun and David, 2025]. Understanding the reasoning behind these choices is essential for materials scientists aiming to optimize synthesis protocols and develop new materials with desirable properties through more efficient processes. Additionally, materials discovery efforts, especially those aided by high-throughput computations [Curtarolo et al., 2013, Kononova et al., 2019], can be accelerated by prioritizing synthesis recipes that are grounded in the most promising mechanistic insights. Per our discussion with materials science researchers, writing the explanations for a synthesis recipe from scratch would take a well-trained materials science PhD student 1-2 hours.
@@ B.3.3 Data Acquisition and Preprocessing
Despite the availability of large-scale synthesis recipe datasets [Kononova et al., 2019, Jain et al., 2013], these curated recipes are not accompanied by explanations. Even for well-trained materials science PhDs, writing explanations for synthesis recipes from scratch is challenging and time-consuming. In contrast, given statements about the synthesis recipe, undergraduate-level chemistry students are able to distinguish between explanations and non-explanations. Therefore, we adopt a semi-automated approach to collect synthesis recipes along with corresponding explanations from existing literature.
@@
We retrieve papers containing the keyword “solid-state synthesis” from Science [^17] https://www.science.org and Wiley [^18] https://onlinelibrary.wiley.com, as the aspects of interest are more relevant to solid-state synthesis. The text of the paper is extracted directly from the HTML version of the paper. During the automated extraction step, we generate five different outputs using GPT-4o for each paper to increase the recall of possible explanations. To improve the faithfulness of the extraction, we also prompt the model to provide the source sentences corresponding to each explanation. The prompt is shown in Tab.˜10, which is based on the rubric designed in Section˜B.3.5. Feedback from the annotators indicates that some explanations extracted by the LLMs describe properties of the synthesized materials (e.g., crystal structure, conductivity) rather than justifying the synthesis steps themselves.
@@ B.3.4 Illustrative Example
An illustrative example is shown in Tab.˜11. All samples for T3 are publicly available.
@@
Additionally, Tab.˜12 presents the model prompt.
@@
{internallinenumbers*} You are a materials science researcher. Given a synthesis recipe that includes the target material, selected precursors, and synthesis steps, your task is to justify the key decisions made in the recipe. This includes explaining the rationale behind the choice of precursors, reaction conditions, and processing steps, using relevant principles such as structural compatibility, chemical reactivity, and desired phase formation. Output the explanation rationales as a list of bullet points, where each bullet point contains complete sentences.
@@ B.3.6 Checklist-mapped Reference
Synthesis explanation generation does not involve the checklist mapping process, as the annotation process directly produces the reference checklist. An example checklist-mapped reference is shown in the illustrative example in Tab.˜11.
@@ B.4.1 Task Definition
Assessing pedagogical alignment involves evaluating how well a large language model (LLM) can replicate effective teaching strategies, such as offering scaffolded guidance without directly revealing the answer [Paul, 1995, Chi, 2009, Hattie and Timperley, 2007, Sonkar et al., 2024]. One possible approach is to evaluate LLMs through real-world interactions with students. However, this poses challenges, as student responses can vary unpredictably across different models [Gao et al., 2025]. Additionally, using the same group of students to test multiple LLMs introduces complications: once students are exposed to a topic in one session, their familiarity with the material can influence their performance in subsequent sessions, making it difficult to isolate the effectiveness of each system. Conversely, using different students for the same set of topics may lead to non-standardized assessments due to variations in their prior knowledge and skill levels. The objective is to assess whether LLMs can provide pedagogically sound feedback that supports student learning without directly giving away answers.
To provide a repeatable and standardized framework for evaluating the pedagogical abilities of different LLMs, we propose a task focused on assessing the effectiveness and pedagogical alignment of feedback within a tutor-student dialogue. Each task instance presents a multi-turn conversation in which a student attempts to solve a biology problem with the guidance of a tutor. The tutor employs a step-by-step problem-solving strategy, breaking the larger problem into smaller sub-problems and promoting active learning by guiding the student through each segment [Chi, 2009]. This process is sequential: the student addresses one sub-problem at a time before moving on to the next. The conversational context includes the full dialogue history up to the current sub-problem, concluding with the student’s latest response.
@@ B.4.2 Task Significance
First, LLMs may produce inaccurate reasoning [Li et al., 2023a] or rely on incorrect information [Razafinirina et al., 2024] when generating feedback, which can hinder rather than support student learning. Second, while LLMs are generally optimized to be helpful and harmless, the operational definition of “helpfulness” may not align with pedagogical goals. Specifically, a fundamental strategy in conceptual learning is to guide students through problem-solving processes without directly revealing the answer. In contrast, LLMs are often trained to prioritize immediate assistance, which may conflict with the educational objective of fostering deep understanding through indirect guidance.
@@ B.4.3 Data Acquisition and Preprocessing
We first provide a background on the source of the data that will be used for building this task. We used the multi-turn student-tutor conversations simulated using GPT-4o by [Sonkar et al., 2023]. These dialogues are grounded in the socio-constructivist model of learning [Stone, 1998], where the tutor adopts a supportive and encouraging tone while offering feedback through Socratic questioning and indirect hints, rather than direct instruction. An illustrative example of such a conversation is provided in Tab.˜15. Each simulated interaction features a student working through a complex biology problem that requires multi-step reasoning and problem decomposition, with the tutor guiding the student through the process, making the dataset well-suited for our evaluation. Human annotators validated each simulated example to ensure factual accuracy and alignment with the socio-constructivist learning model.
@@
In contrast, we concentrate solely on the non-affective components of the feedback, as prior studies have shown minimal influence of positive language and praise on student outcomes [Kluger and DeNisi, 1996, Ferris, 1997]. Our primary objective is to develop evaluation criteria that assess the effectiveness of feedback in promoting conceptual understanding. Studies have consistently found that feedback elements which identify problems and errors and propose solutions are the most beneficial for student learning [Hayes et al., 1987, Matsumura et al., 2002, Bitchener et al., 2005, Sugita, 2006]. Accordingly, our task is designed to evaluate an LLM’s ability to generate feedback that provides accurate problem diagnosis and rectification. While indirect feedback—such as hints or questions that avoid explicitly identifying the mistake or providing the answer—can foster meta-cognitive and conceptual learning, its effectiveness may diminish when students repeatedly struggle with the same sub-problem [Westmacott, 2017]. Accordingly, our evaluation also emphasizes the LLM’s ability to generate contextually appropriate feedback based on the nature of the student’s mistakes for a given sub-problem. In cases where the student makes repeated errors on the same sub-problem, the LLM should accurately recognize this pattern and respond with explicit error identification and corrective guidance. In this section, we explain how to get fine-grained information from the seed data collected by [Sonkar et al., 2023].
@@
Consider a single response of the tutor from the example shown in Tab.˜15. We focus on the second response from the tutor and provide fine-grained information associated with it in Tab.˜13.
@@
To assess effectiveness in error identification and rectification, we extract these components from the feedback generated by the simulated tutor. This extraction is carried out using GPT-4o for automated annotation, with subsequent human verification to ensure the accuracy of the identified issues and corresponding corrective suggestions. Since the task does not require deep domain-specific knowledge, the manual verification was performed by one of the authors with strong proficiency in English. An illustrative example of the extracted elements from a single data point is provided in Tab.˜14.